# Exploring End-to-end Differentiable Neural Charged Particle Tracking – A Loss Landscape Perspective

### v1

@@ 5.3 Representational and Functional Similarities
As illustrated in Figure 7, both PTT and PAT-based models show substantial functional disagreement in a similar range, with median disagreement rates around 15%. We want to specifically point out the increase in disagreement for models with higher interpolation values (see Appendix F for the statistical significance analysis), suggesting choosing smaller $\lambda$ values for improved stability of the model. Further, the experiments demonstrate, similar to the representational similarity results, a significant increase in disagreement for combinations containing both PAT and PTT variants. Here, we can observe median disagreement rates of approximately 20%, corresponding to an increase in prediction instability of approximately 33%.
5.4 Global Connectivity of Trained Models
@@
Hamming loss [×$\times 1e^{-3}$] ($\downarrow$) True positive rate [%] ($\uparrow$)
@@
PAT ($\lambda=25$) Linear 6.64$\pm$0.09 228.58$\pm$110.62 37.73$\pm$19.92 75.01$\pm$11.77 99.32$\pm$0.01 95.80$\pm$2.16
PAT ($\lambda=25$) Bézier (Ham.) 6.64$\pm$0.07 8.14$\pm$0.53 7.11$\pm$0.10 99.12$\pm$0.09 99.32$\pm$0.01 99.26$\pm$0.01
PAT ($\lambda=25$) Bézier (BCE) 6.65$\pm$0.11 7.78$\pm$0.13 7.35$\pm$0.06 99.18$\pm$0.02 99.32$\pm$0.01 99.24$\pm$0.01
PAT ($\lambda=50$) Linear 6.47$\pm$0.12 251.09$\pm$102.46 43.46$\pm$22.46 71.57$\pm$10.24 99.34$\pm$0.01 95.01$\pm$2.39
PAT ($\lambda=50$) Bézier (Ham.) 6.46$\pm$0.12 8.01$\pm$0.40 7.00$\pm$0.09 99.15$\pm$0.05 99.34$\pm$0.01 99.28$\pm$0.01
PAT ($\lambda=50$) Bézier (BCE) 6.47$\pm$0.08 7.76$\pm$0.12 7.35$\pm$0.05 99.18$\pm$0.02 99.34$\pm$0.01 99.24$\pm$0.01
PAT ($\lambda=75$) Linear 6.49$\pm$0.04 345.56$\pm$133.60 82.49$\pm$44.28 63.04$\pm$13.58 99.34$\pm$0.00 91.13$\pm$4.60
PAT ($\lambda=75$) Bézier (Ham.) 6.47$\pm$0.05 8.22$\pm$0.79 7.02$\pm$0.16 99.11$\pm$0.12 99.34$\pm$0.00 99.27$\pm$0.02
PAT ($\lambda=75$) Bézier (BCE) 6.51$\pm$0.04 8.12$\pm$0.72 7.40$\pm$0.08 99.16$\pm$0.05 99.34$\pm$0.00 99.23$\pm$0.01
PTT Linear 6.70$\pm$0.07 332.39$\pm$116.90 107.24$\pm$42.20 63.67$\pm$11.94 99.32$\pm$0.01 87.59$\pm$4.31
PTT Bézier (Ham.) 6.70$\pm$0.07 15.33$\pm$5.54 9.56$\pm$1.07 97.98$\pm$0.94 99.32$\pm$0.01 98.91$\pm$0.18
PTT Bézier (BCE) 6.69$\pm$0.07 7.60$\pm$0.21 7.07$\pm$0.06 99.21$\pm$0.03 99.32$\pm$0.01 99.28$\pm$0.01
@@
Hamming loss [×$\times 1e^{-3}$] ($\downarrow$) True positive rate [%] ($\uparrow$)
@@
$\text{GNN}_{PTT}\rightarrow\text{GNN}_{PAT}$ Linear 6.60$\pm$0.13 337.23$\pm$169.20 117.42$\pm$76.26 63.90$\pm$17.66 99.33$\pm$0.01 87.12$\pm$8.06
$\text{GNN}_{PTT}\rightarrow\text{GNN}_{PAT}$ Bézier (Ham.) 6.60$\pm$0.13 11.32$\pm$3.25 8.06$\pm$0.58 98.65$\pm$0.52 99.33$\pm$0.01 99.13$\pm$0.08
$\text{GNN}_{PTT}\rightarrow\text{GNN}_{PAT}$ Bézier (BCE) 6.59$\pm$0.12 7.75$\pm$0.36 7.19$\pm$0.09 99.19$\pm$0.03 99.33$\pm$0.01 99.26$\pm$0.01
Given the findings of Section 5.3, demonstrating substantial disagreements in model predictions, we now aim to quantify the global shape or connectivity of the loss landscape. Table 2 and Table 3 summarize the mode connectivity, comparing connecting Bézier curves between model parameters as specified in Section 5.2. To get a more profound understanding of the global landscape for prediction and task loss, we generate curves using both the BCE-loss on the predicted edge-scores and the Hamming loss on the assigned edges. Here, we take advantage of our findings in Section 5.2.1, suggesting a strong agreement between Hamming-loss and BCE-loss for low-loss values. In addition, we provide the linear connectivity between two models as a baseline. Further, Neyshabur et al. (2020); Juneja et al. (2022) and Lubana et al. (2023) demonstrate in their work that trained models deficient in linear connectivity, are mechanistic dissimilar [^7] Trained models using different mechanisms or representations for predictions.. This allows us to gain further insides on the stability of the trained models from a loss landscape perspective, supporting the results in Section 5.3.
@@
We find similar results demonstrating strong nonlinear connectivity between trained models of PAT and PTT mixtures, indicating a good global connectivity of the modes without noticeable loss barriers that could indicate more pronounced differences in the optimization process (see Table 3). However, the high loss values of the linear connectivity ($\text{mean}_{PTT\rightarrow PAT}$: 117.42$\pm$76.26, vs. e.g., $\text{mean}_{PTT}$: 107.24$\pm$42.2) demonstrate a noticeable impact of mixing the training paradigms on the mechanistic similarity. In addition, we find similar to the connecting curves of PTT minimizing the Hamming-loss noticeable higher loss values compared to the BCE-loss.
@@ 6 Discussion and Conclusion
In this work, we propose and explore E2E differentiable neural charged particle tracking using graph neural networks. We integrate combinatorial optimization mechanisms used for generating disconnected track candidates from predicted edge scores directly into the training pipeline by leveraging mechanisms from decision-focused learning. By optimizing the whole track reconstruction pipeline in an E2E fashion, we obtain comparable results to two-step training, minimizing the BCE-loss of raw edge scores. However, we demonstrate the predict-and-track approach can provide gradient information that can be propagated throughout the reconstruction process. While the usage for simply optimizing the reconstruction performance of the network is limited, providing gradient information is highly valuable for various use-cases such as uncertainty propagation, reconstruction pipeline optimization and allows the construction of complex, multistep architectures, potentially allowing to reduce the combinatorial nature of particle tracking by ensuring unique hit assignment.
@@
Our analysis demonstrates, beside the similar reconstruction results obtained for different interpolation values and the general understanding that the exact choice of the hyperparameter is insignificant (Vlastelica et al., 2020), a coupling of the choice of lambda on the prediction stability of the network. Selecting lower values for lambda is thus especially important in critical applications where a consistency of outputs is desirable or even essential.
Moving forward, we plan to expand the framework by incorporating auxiliary losses for downstream tasks, especially for the use-case of pCT. Additionally, we will continue extending our studies to explore other potential combinatorial solvers for generating disconnected track candidates from predicted edge scores, as well as investigate additional network architectures to gain a more profound understanding of the impact of end-to-end optimization on charged particle tracking.
@@
Max Aehlea, Johan Almeb, Gergely Gábor Barnaföldic, Tea Bodovab, Vyacheslav Borshchovd, Anthony van den Brinke, Mamdouh Chaarb, Viljar Eikelandb, Gregory Feofilovf, Christoph Garthg, Nicolas R. Gaugera, Georgi Genovb, Ola Grøttvikb, Håvard Helstruph, Sergey Igolkinf, Ralf Keideli, Chinorat Kobdajj, Tobias Kortusa, Viktor Leonhardtg, Shruti Mehendaleb, Raju Ningappa Mulawadei, Odd Harald Odlandk, b, George O’Neillb, Gábor Pappl, Thomas Peitzmanne, Helge Egil Seime Pettersenk, Pierluigi Piersimonib,m, Maksym Protsenkod, Max Rauchb, Attiq Ur Rehmanb, Matthias Richtern, Dieter Röhrichb, Joshua Santanai, Alexander Schillingi, Joao Secoo, p, Arnon Songmoolnakb, j, Ákos Sudárc, q, Jarle Rambo Sølier, Ganesh Tambaves, Ihor Tymchukd, Kjetil Ullalandb, Monika Varga-Kofaragoc, Boris Wagnerb, RenZheng Xiaob, v, Shiming Yangb, Hiroki Yokoyamae,
@@ Acknowledgments
This work was supported by the German federal state Rhineland-Palatinate (Forschungskolleg SIVERT) and by the Research Council of Norway (Norges forskningsråd), the German National High-Performance Computing (NHR) association for the Center NHR South-West and the University of Bergen, grant number 250858. The simulations and training were partly executed on the high-performance cluster "Elwetritsch" at the University of Kaiserslautern-Landau, which is part of the "Alliance of High-Performance Computing Rhineland-Palatinate" (AHRP). We kindly acknowledge the support of the regional university computing center (RHRK). Tobias Kortus and Nicolas Gauger gratefully acknowledge the funding of the German National High-Performance Computing (NHR) association for the Center NHR South-West. The ALPIDE chip was developed by the ALICE collaboration at CERN.
@@ Appendix A Edge Filter Results
In Section 4.1 we mentioned the necessity of performing edge filtering, minimizing the size of both the hit- and line graph representations, improving reconstruction speeds and minimizing memory requirements during training and inference. Selecting sufficient thresholds, given the stochastic nature of particle scattering in material, requires a trade-off between reduced graph size and the number of illegitimately removed true edges. Figure 8 shows the fraction of removed number of edges in the graph opposed to the fraction of removed true edges, generated for 100 graphs of the train dataset selected uniformly over all phantom thicknesses. Based on the results in Figure 8 we select the thresholds $\theta_{d}=400$ mrad for the edges in the calorimeter layer and $\theta_{t}=200$ mrad for all edges contained in the tracking layer.
@@ Appendix C Reconstructed Particle Tracks
Figure 9: Reconstructed particle tracks generated for multiple readout frames with 100mm water phantom and 100 $p^{+}/F$ using PAT ($\lambda=25$) model. Marked in green are correctly reconstructed track segment, marked in red are incorrectly reconstructed track segments and marked in orange are correct reconstructions (same particle ID) following the wrong primary particle.
Appendix D Implementation Details for Loss Landscape Evaluations in Section 5.2
@@
We closely follow the implementation details for generating connecting, curves provided by (Garipov et al., 2018). However, instead of updating the batch normalization parameters for every t, we follow a similar approach used for generating the loss landscapes Yang et al. (2021). We use constant parameters determined by the first curve anchor, providing us with reasonable normalization values. Each curve is optimized using 1000 iterations with a batch size of 8 graphs. We found that increasing the training iterations resulted with worse connectivity results for PTT-based combinations, minimizing the hamming loss. We use the same parallelization strategy as used for calculating representational and functional similarities.
Appendix E Reproducibility

### v2

@@ 5.3 Representational and Functional Similarities
As illustrated in Figure 7, both PTT and PAT-based models show substantial functional disagreement in a similar range, with median disagreement rates around 15%. We want to specifically point out the increase in disagreement for models with higher interpolation values (see Appendix G for the statistical significance analysis), suggesting choosing smaller $\lambda$ values for improved stability of the model. Further, the experiments demonstrate, similar to the representational similarity results, a significant increase in disagreement for combinations containing both PAT and PTT variants. Here, we can observe median disagreement rates of approximately 20%, corresponding to an increase in prediction instability of approximately 33%.
Interpretation:
We find that the substantial degree of prediction instability and different levels of similarity in the learned representations pose a considerable risk, potentially influencing the overall performance, effectiveness, and reliability of entire component-based processing pipelines (see Section 3.2) in the application domain. While we do not expect significant effects on, e.g., the average reconstruction results of pCT, especially for homogeneous materials, the prediction instability could have a measurable impact on the contrast and spatial resolution.The instability in downstream reconstruction remains unquantified and is left for future work to maintain the conciseness of this initial study. By propagating information from a downstream task, E2E optimization provides an important tool to restrict the solution set and thus control the quality of learned solutions. While some of the instabilities can be avoided to improve reproducibility, other sources of randomness introduced by processing hardware and implementation, e.g., nondeterministic scatter operation of PyTorch Geometric (Klabunde & Lemmerich, 2023) still remain and display a noteworthy source of uncertainty and unreliability, propagated throughout the entire processing pipeline.
5.4 Global Connectivity of Local Minima
Given the findings of Section 5.3, demonstrating substantial disagreements in model predictions, we now aim to quantify the global shape and connectivity of the loss landscape. Table 2 and Table 3 summarize the mode connectivity, comparing connecting Bézier curves between model parameters as specified in Section 5.2. To get a more profound understanding of the global landscape for prediction and task loss, we generate curves using both the BCE-loss on the predicted edge-scores and the Hamming loss on the assigned edges. Here, we take advantage of our findings in Section 5.2.1, suggesting a strong agreement between Hamming-loss and BCE-loss for low-loss values. In addition, we provide the linear connectivity between two models as a baseline. Further, Neyshabur et al. (2020); Juneja et al. (2022) and Lubana et al. (2023) demonstrate in their work that trained models deficient in linear connectivity, are mechanistic dissimilar [^9] Trained models using different mechanisms or representations for predictions.. This allows us to gain further insides on the stability of the trained models from a loss landscape perspective, supporting the results in Section 5.3.
@@ 5.4 Global Connectivity of Local Minima
Hamming loss $\times 1e^{-3}$] ($\downarrow$) True positive rate [%] ($\uparrow$)
@@
[…]
@@
Hamming loss $\times 1e^{-3}$] ($\downarrow$) True positive rate [%] ($\uparrow$)
@@
[…]
@@
[…]
Interpretation:
While analyzing the global connectivity of loss landscape does not replace the analysis of E2E-differentiability in an entire processing pipeline, we argue that good global connectivity of local loss basins which are interconnected via a low-loss non-linear path is a good initial indication that E2E particle tracking can be efficiently augmented by gradients originated from a downstream reconstruction or analysis loss. Finding a parameter configuration in the low-loss region that minimizes the performance of the downstream task is thus likely effective without getting stuck in local regions, that are separated by high loss barriers. Further, the low mechanistic, representational and functional similarities (see Section 5.3 and 5.4) suggests that a diverse set of reconstruction policies can be reached in those interconnected regions.
@@ 6 Discussion and Conclusion
In this work, we propose and present first studies on E2E differentiable neural charged particle tracking using graph neural networks. We integrate combinatorial optimization mechanisms used for generating disconnected track candidates from predicted edge scores directly into the training pipeline by leveraging mechanisms from decision-focused learning. By optimizing the whole track reconstruction pipeline in an E2E fashion, we obtain comparable results to two-step training, minimizing the BCE-loss of raw edge scores. However, we demonstrate the predict-and-track approach can provide gradient information that can be propagated throughout the reconstruction process. While the usage for simply optimizing the reconstruction performance of the network is limited, providing gradient information is highly valuable for various use-cases such as uncertainty propagation, reconstruction pipeline optimization and allows the construction of complex, multistep architectures, potentially allowing to reduce the combinatorial nature of particle tracking by ensuring unique hit assignment.
@@
Our analysis demonstrates, beside the similar reconstruction results obtained for different interpolation values and the general understanding that the exact choice of the hyperparameter is non-critical (Vlastelica et al., 2020), a coupling of the choice of $\lambda$ on the prediction stability of the network and an inverse dependency on the convergence speed. Thus, selecting large interpolation constants, improves the informativeness of gradient information, resulting in faster convergence. The gain in convergence speed, however, comes at the cost of decreased robustness. Selecting lower values for lambda is thus especially important in critical applications where a consistency of outputs is desirable or even essential.
Moving forward, we plan to expand the framework by integrating E2E tracking into reconstruction code incorporating auxiliary losses for downstream tasks, especially for the use-case of pCT. Additionally, we will continue extending our studies to explore the out-of-distribution robustness and general robustness to distribution shifts introduced by simulation to reality gaps. Further, we plan extending our work to other potential combinatorial solvers for generating disconnected track candidates from predicted edge scores, as well as investigating additional network architectures to gain a more profound understanding of the impact of end-to-end optimization on charged particle tracking.
@@
Max Aehlea, Johan Almeb, Gergely Gábor Barnaföldic, Tea Bodovab, Vyacheslav Borshchovd, Anthony van den Brinke, Mamdouh Chaarb, Viljar Eikelandb, Gregory Feofilovf, Christoph Garthg, Nicolas R. Gaugera, Georgi Genovb, Ola Grøttvikb, Håvard Helstruph, Sergey Igolkinf, Ralf Keideli, Chinorat Kobdajj, Tobias Kortusa, Viktor Leonhardtg, Shruti Mehendaleb, Raju Ningappa Mulawadei, Odd Harald Odlandk, b, George O’Neillb, Gábor Pappl, Thomas Peitzmanne, Helge Egil Seime Pettersenk, Pierluigi Piersimonib,m, Maksym Protsenkod, Max Rauchb, Attiq Ur Rehmanb, Matthias Richtern, Dieter Röhrichb, Joshua Santanai, Alexander Schillingi, Joao Secoo, p, Arnon Songmoolnakb, j, Ákos Sudárc, q, Jarle Rambo Sølier, Ganesh Tambaves, Ihor Tymchukd, Kjetil Ullalandb, Monika Varga-Kofaragoc, Boris Wagnerb, RenZheng Xiaob, v, Shiming Yangb, Hiroki Yokoyamae
@@ Acknowledgments
We thank the anonymous reviewers for their thoughtful comments that greatly improved the final version of this manuscript. This work was supported by the German federal state Rhineland-Palatinate (Forschungskolleg SIVERT), the Research Council of Norway (Norges forskningsråd), and the University of Bergen, grant number 250858. The simulations and training were partly executed on the high-performance cluster "Elwetritsch" at the University of Kaiserslautern-Landau, which is part of the "Alliance of High-Performance Computing Rhineland-Palatinate" (AHRP). We kindly acknowledge the support of the regional university computing center (RHRK). Tobias Kortus and Nicolas Gauger gratefully acknowledge the funding of the German National High-Performance Computing (NHR) association for the Center NHR South-West. The ALPIDE chip was developed by the ALICE collaboration at CERN.
@@ Appendix A Edge Filter Results
In Section 4.1 we mentioned the necessity of performing edge filtering, minimizing the size of both the hit- and line graph representations, improving reconstruction speeds and minimizing memory requirements during training and inference. Selecting sufficient thresholds, given the stochastic nature of particle scattering in material, requires a trade-off between reduced graph size and the number of illegitimately removed true edges. Figure 8 shows the fraction of removed number of edges in the graph opposed to the fraction of removed true edges, generated for 100 graphs of the train dataset selected uniformly over all phantom thicknesses. Based on the results in Figure 8 we select the thresholds $\theta_{d}=400$ mrad for the edges in the calorimeter layer and $\theta_{t}=200$ mrad for all edges contained in the tracking layer, significantly reducing the number of total edges, while limiting the impact on the final reconstruction quality to a minimum. Error rates for both $\theta_{t}$ and $\theta_{d}$ are presented in Figure 8.
@@ Appendix C Reconstructed Particle Tracks
Figure 9: Reconstructed particle tracks generated for multiple readout frames with 100mm water phantom and 100 $p^{+}/F$ using PAT ($\lambda=25$) model. Marked in green are correctly reconstructed track segments, marked in red are incorrectly reconstructed track segments and marked in orange are correct reconstructions (same particle ID) following the wrong primary particle.
Appendix D Ablation Study: Hit-graph and Line-graph Representations
While demonstrating good training performance, utilizing a line-graph representation for particle tracking comes at the cost of increased graph size and thus computational overhead over the original hit-graph representation. To verify the necessity and usefulness of this transformation over a default hit graph parameterization, this section provides additional ablation results comparing hit- and line-graph representations in terms of convergence behavior, reconstruction performance and inference speed.
Reference results utilizing a hit-graphs are generated for all model configurations outlined in Section 5 by replacing the modified architecture with the original interaction-network implementation (Battaglia et al., 2016; 2018) used in DeZoort et al. (2021). The assignment operation, described in Section 4.2 is left as is. We parameterize both vertex- ($\bm{v}_{i}$) and edge features ($\bm{e}_{ij}$) following previous work by (Kortus et al., 2023), using energy deposition and global position ($\Delta E,x,y,z$) and a description of the edge in terms of spherical coordinates ($r,\theta,\phi$) orthogonal to the detector plane respectively (see Figure 2). Table 5 summarizes the average node and edge count ($\pm$ standard deviation) for each evaluated configuration. For every interaction-network configuration operating on a hit-graph, the same hyperparameters from Appendix B are utilized except for the hidden dimension $d_{hidden}$, which is increased to 64. While we found that the interaction network operating on the hit-graph benefits from the increased hidden dimensionality, we find similar to DeZoort et al. (2021) that the model saturates around the selected value and further increase in parameter count does not yield any noteworthy improvement.
Table 5: Average node- and edge count of hit- and line-graph representations ($\pm$ standard deviation), for simulated data with 100, 150 and 200 m water phantoms and 5, 100 and 150 primaries per frame (p+/F).
100 mm Water Phantom 150 mm Water Phantom 200 mm Water Phantom
p+/F Repr. #nodes #edges #nodes #edges #nodes #edges
[…]
Figure 10: True positive and false positive rate of assignments, as well as purity and efficiency of track reconstruction, evaluated for PAT and PTT with hit-graph representation on validation set as a function of optimisation steps.
Table 6: Reconstruction performance for hit- and line-graph representations, measured as purity $p$ and efficiency $\epsilon$ for water phantoms of 100, 150 and 200 mm thickness and 50, 100 and 150 primaries per frame $p^{+}/F$. The results for $\mathcal{G}_{H}$ were previously presented in Table 1.
[…]
Training performance:
Figure 10 visualizes analogous to Figure 4 the performance metrics on training and validation set as a function of training steps. We find that restricting the readout representation to a hit-graph severely impedes training performance. Across all model configurations, we find a significant drop across all measured quantities. Especially, end-to-end optimization strategies suffer from the reduced information density encoded in the line-graph representation. This results both in decreased convergence speed and stability, as well as reduced reconstruction quality.
Reconstruction performance:
Table 6 further confirms the previous finding, comparing the reconstruction quality of the hit-graph baseline ($\mathcal{G}_{H}$) with the line graph ($\mathcal{G}_{L}$) results, first presented in Table 1. We find a substantial drop in reconstruction purity and efficiency across all tested phantom geometries and particle densities when the line graph transformation is omitted. This effect is especially pronounced for the end-to-end trained baseline, confirming the necessity of the line-graph transformation and the importance of the direct parameterization of segment-pair features, capturing the scattering behavior of particles, for both two-stage and end-to-end approaches. In line with (Kortus et al., 2023), highlighting the importance of positional encodings for segment pairs, we specifically credit the significant gain in performance to the inductive bias, introduced by the encodings, simplifying learning the underlying interaction mechanisms defined by multiple Coulomb scattering (see. Section 3.1). This effect further enables the use of smaller models, enhancing parameter efficiency compared to networks that operate on the hit-graph parameterization.
Computational complexity:
Figure 11 outlines the estimated floating-point operations (FLOP) of graph and line-graph interaction networks, operating on various phantom and density configurations [^10] All FLOP estimates were calculated using the FlopCountAnalysis tool provided as a part of fvcore (FAIR, 2019).. We provide averaged results obtained over all readout frames for each given configuration.
Figure 11: Estimated number of floating-point operations (FLOP) for hit-graph and line-graph interaction networks, evaluated for readout frames of various phantom geometries and particle densities.
Ultimately, Figure 12 presents and compares the average reconstruction times of readout frames for both hit-graph and line-graph representation to quantify the runtime penalty of line-graph representations. Both runtimes for the interaction-network prediction (including the line-graph transform) and total reconstruction are presented for various water phantom geometries and particle densities. All results are generated on an NVIDIA A100 GPU for all readout frames in the test set, with 10 evaluations for each readout frame.
Figure 12: Average runtime of edge-classification charged particle tracking (predictive stage and full reconstruction.) operating on hit-graph ($\mathcal{G}_{H}$) and line-graph ($\mathcal{G}_{L}$) representation for readout frames of various water phantom geometries and particle densities.
We find across all configurations that the line-graph interaction-network noticeably impedes the runtime of the predictive phase, both in terms of FLOP and runtimes, due to the significantly increased and denser connectivity of the graph (see Table 5). Yet, the overall impact on the total reconstruction runtime is manageable for the defined scope, outweighing the increase in runtime by the significant gain in reconstruction quality.
Appendix E Implementation Details for Loss Landscape Evaluations in Section 5.2
@@ Appendix D Ablation Study: Hit-graph and Line-graph Representations
We closely follow the implementation for generating connecting, curves provided by (Garipov et al., 2018). However, instead of updating the batch normalization parameters for every t, we follow a similar approach used for generating the loss landscapes in Yang et al. (2021). We use constant parameters determined by the first curve anchor, providing us with reasonable normalization values. Each curve is optimized using 1000 iterations with a batch size of 8 graphs. We found that increasing the training iterations resulted with worse connectivity results for PTT-based combinations, minimizing the hamming loss. We use the same parallelization strategy as used for calculating representational and functional similarities.
Appendix F Reproducibility